7(3x + 4) = 8(2x + 5) + 13

2(x - 3) - 3(x + 5)= 3(x - 2) - 7

3x - 5(3x - 7)= 2(x + 9) + 45

Answer by checkley71(8403)

7(3X+4)=8(2X+5)+13
21X+28=16X+40+13
21X-16X=53-28
5X=25
X=25/5
X=5 ANSWER.
PROOF
7(15+4)=8(10+5)+13
7*19=8*15+13
133=120+13
133=133
----------------------------------------
HERE'S THE SOLUTION TO THE FIRST PROBLEM.
USING THE SAME PROCEDURE YOU SHOULD BE ABLE TO SOLVE THE OTHER TWO.
WE PROVIDE SOLUTION.
WE ARE NOT HERE TO DO HOMEWORK.